• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name weinstein

    Log entry time 16:01:31 on March24,2005

    Entry number 142329

    keyword=shift summary

    run 2902
    Bigbite VME crate not responding, ask for access
    continue run (since at least we're getting neutron data)
    0830 access to power cycle ROC5
    run 2903 to test ROC5
    0839 looks like it worked
    now MCC does harp scans
    some confusion about where the harp HV controls are, can't reach PD
    0926 request real beam
    0927 HRS-L Q3 drifting a little (for about a minute or so)
    0928 run 2904 started
    0936 ROC5 not reporting, end run at 23K events
    on Hall A, slow controls screen, reset SRC Crates: Trigger supervisor and SFI fastbus
    2905 is junk
    0942 run 2906 started. JUNK
    neutron scalers have many zeros. lots of HV channels off in BB and n array. beam position was way off. reset HV and ask for beam centered
    10:00 run 2907 started
    10:27 run 2907 ended with 150k ev (MCC 'looking at the beam' now)
    10:35 MCC is done, beam to be restored
    start run 2908
    10:48 end run 2908 w/69k ev in order to tune through the Compton chicane
    12:26 physics beam back
    12:26-13:02 run 2909 200k events
    13:20 Right VDC HV is low (3.46/3.73) but VDC efficiencies looked fine on the replay
    13:20-13:48 run 2910 200k events
    T1 (right arm) is running at 2500 Hz and ps1 = 50. This is too much and runs are too short. Double ps1 to 100.
    14:46 end run 2911 200k events
    14:50 lose beam (access to north linac to find vacuum problem)
    14:50 end run 2912 26k events (mostly no beam)